Find the potential o when F=-V¢, a conservative force of field defined by + = (3x’yz – 3y)i + (x?z – 3x)j + (x*y+ 2zł
Answer #1

I have done it for you in detail. Kindly go through.

We need to find such that F=-00 where F = (x+yz - 3y) 7 + (vºz – 3x)] +(y+22)ā. + F-ф — Ф. зха? - За 3 opy х2 - 3x 2 ф, х hosNow o becomes &= 3xy - xyz + g(2) Now differentiate it with respect to z. we get, 21 -zy + g(x). ~ -xy-22 = -xej tg(a) (by